# DRCT: Saving Image Super-Resolution away from Information Bottleneck

### v2

@@ Abstract
In recent years, Vision Transformer-based applications to low-level vision tasks have achieved widespread success. Unlike CNN-based models, Transformers are more adept at capturing long-range dependencies, enabling the reconstruction of images utilizing information from non-local areas. In the domain of super-resolution, Swin-transformer-based approaches have become mainstream due to their capacity to capture global spatial information and their shifting-window attention mechanism that facilitates the interchange of information between different windows. Many researchers have enhanced image quality and network efficiency by expanding the receptive field or designing complex networks, yielding commendable results. However, we observed that spatial information tends to diminish during the forward propagation process due to increased depth, leading to a loss of spatial information and, consequently, limiting the model’s potential. To address this, we propose the Dense-residual-connected Transformer (DRCT), aimed at mitigating the loss of spatial information through dense-residual connections between layers, thereby unleashing the model’s potential and enhancing performance. Experiment results indicate that our approach is not only straightforward but also achieves remarkable efficiency, surpassing state-of-the-art methods and performing commendably at NTIRE2024.
@@ 1 Introduction
Figure 1: The feature map intensity on various benchmark datasets. We observed that feature map intensities decrease sharply at deeper network levels, indicating potential information loss. We propose DRCT, stabilizes this by enhancing receptive fields and adding dense-connections within residual blocks to mitigate information bottlenecks, thereby improving performance with a simpler model design.
@@
Following the foundational studies, CNN-based strategies [10, 25, 8, 56, 32, 31] have predominantly governed the super-resolution domain for an extended period. These strategies largely leverage techniques such as residual learning [53, 43, 38, 19, 26], or channel attention [46, 52] for developing network architectures, significantly propelling the progress of super-resolution models forward.
CNN-based networks have achieved notable success in terms of performance. However, their inherent limitations stem from the parameter-dependent scaling of the receptive field and the local nature of convolutional interactions, which are independent of content. These characteristics limit CNNs’ ability to effectively model long-range dependencies.
To overcome the limitations associated with CNN-based networks, researchers have introduced Transformer-based SISR networks that leverage the capability to model long-range dependencies, thereby enhancing SISR performance. Notable examples include IPT [3] and EDT [22], which utilize pre-training on the ImageNet [9] dataset to fully leverage Transformer capabilities for superior SISR results. SwinIR [24], inspired by the Swin-transformer [27] architecture, marks a significant advancement in SISR performance.
@@
While using HAT and SwinIR for inference across various datasets, we observed a common phenomenon: the intensity distribution of the feature maps undergoes more substantial changes as the network depth increases. This indicates the spatial information and attention intensity learned by the model. However, there’s often a sharp decrease towards the end of the network, shrinking to a smaller range. This intriguing pattern suggests that such abrupt changes might be accompanied by a loss of information, indicating the presence of an information bottleneck.
Inspired by a series of works by Wang et al., such as the YOLO-family [42, 39], CSPNet [40], and ELAN [41], we consider that network architectures based on SwinIR, despite significantly enlarging the receptive field through shift-window attention to address the issue of insufficient receptive fields in CNNs, are prone to gradient bottlenecks due to the loss of spatial information as network depth increases. This implicitly constrains the model’s performance.
To address the issue of spatial information loss due to an increased number of network layers, we introduce the Dense-residual-connected Transformer (DRCT), designed to stabilize the forward-propagation process and prevent information bottlenecks. This is achieved by the proposed Swin-Dense-Residual-Connected Block (SDRCB), which incorporates Swin Transformer Layers and transition layers into each Residual Deep Feature Extraction group (RDFEG). Consequently, this approach enhances the receptive field with fewer parameters and a simplified model structure, thereby resulting in improved performance.
@@ 2 Related works
Figure 2: The feature map visualization displays, from top to bottom, SwinIR [24], HAT [5], and DRCT, with positions further to the right representing deeper layers within the network. For both SwinIR and HAT, the intensity of the feature maps is significant in the shallower layers but diminishes towards the network’s end. We consider this phenomenon may imply the loss of spatial information, leading to the limitation and information bottleneck with SISR tasks.
Figure 3: The G-index we propose is calculated by summing the absolute values of $\Delta G_{min}$ and $\Delta G_{max}$, where $\Delta G$ represents the change in feature map intensity between two consecutive layers. This change is measured by the difference in intensity levels, capturing both the minimum and maximum shifts across layers.
@@ 2.1 Vision Transformer-based Super-resolution
IPT [3], a versatile model utilizing the Transformer encoder-decoder architecture, has shown efficacy in super-resolution, along with denoising and deraining tasks. SwinIR [24], building on the Swin Transformer [27] encoder, employs self-attention within local windows during feature extraction, resulting in superior performance. UFormer [44] introduces an innovative local-enhancement window Transformer block, greatly reducing computational demands for processing high-resolution features. Furthermore, it introduces a learnable multi-scale restoration modulator within the decoder to enhance the model’s ability to detect both local and global patterns necessary for image restoration. ART [58] incorporates an attention retractable module to expand its receptive field, thereby enhancing SISR performance. CAT [6] leverages rectangle-window self-attention for feature aggregation, achieving a broader receptive field. It also integrates a locality complementary module to effectively merge global and local information, improving image restoration outcomes. HAT [5] integrates channel attention with window-based self-attention, leveraging both global information and precise local fitting abilities. Additionally, it introduces an overlapping cross-attention module, specifically designed to improve the aggregation of information across different windows, thereby enhancing feature interaction and setting new benchmarks in the field.
@@ 2.2 Auxiliary Supervision and Feature Fusion
Auxiliary Supervision. Deep supervision is a commonly used auxiliary supervision method [20, 33] that involves training by adding prediction layers at the intermediate levels of the model [39, 40, 41]. This approach is particularly prevalent in architectures based on Transformers that incorporate multi-layer decoders. Another popular auxiliary supervision technique involves guiding the feature maps produced by the intermediate layers with relevant metadata to ensure they possess attributes beneficial to the target task [14, 15, 43, 53, 18]. Choosing the appropriate auxiliary supervision mechanism can accelerate the model’s convergence speed, while also enhancing its robustness and performance.
Feature Fusion. Various studies have explored the integration of features across different dimensions in multiple visual tasks [26, 47] to enhance performance. In CNNs, attention mechanisms have been applied to both spatial and channel dimensions to improve feature representation; examples of which include RTCS [16] and SwinFusion [28]. In ViT [11], spatial self-attention is used to model the long-range dependencies between pixels. Additionally, some researchers have investigated the incorporation of channel attention within Transformers [55, 4] to effectively amalgamate spatial and channel information. This integration significantly enhances the modeling capabilities of Transformers.
@@ 3 Problem Statement
Figure 4: The overall architecture of the proposed DRCT and the structure of RDFEG and SDRCB.
@@ 3.1 Information Bottleneck Principle
$\small I(X,X)\geq I(Y,X)\geq I(Y,f_{\theta}(X))\geq I(X,g_{\phi}(f_{\theta}(X)))\geq\ldots\geq I(Y,\hat{Y})$ (1)
@@
In deep neural networks, $f_{\theta}(\cdot)$ and $g_{\phi}(\cdot)$ respectively represent the two consecutive layers in neural network. From equation 1, we can predict that as the number of network layer becomes deeper, the original data will be more likely to be lost.
In term of SISR tasks, the general goal is to find the mapping function $F$ to maximize the mutual information between HR and SR image.
$F(I_{LR};\theta)=I_{SR};\\ \max_{\theta}I(I_{HR};F(I_{LR};\theta))$ (2)
@@ 3.2 Spatial Information Vanish in Super-resolution
Generally speaking, SISR methods [5, 6, 24, 50, 58, 7, 57, 21] can generally divided into three parts: (1) shallow feature extraction, (2) deep feature extraction, (3) image reconstruction.
In these methods, there is almost no difference between shallow feature extraction and image reconstruction. The former is composed of simple convolutional layers, while the latter consists of convolutional layers and upsampling layers. However, deep feature extraction differs significantly. Yet, their commonality lies in being composed of various residual blocks, which can be simply defined as:
$X^{l+1}=X^{l}+f_{\theta}^{l+1}(X^{l})$ (3)
@@
Using residual learning allows the model to only update the differences between layers, rather than predicting the total output of a layer directly. This reduces the difficulty of model training and prevents gradient vanishing. However, according to our observations, while this design effectively transmits spatial information between different residual blocks, there may still be information loss.
Because the information within a residual block may not necessarily maintain spatial information, this ultimately leads to non-smoothness in terms of feature map intensity (refer to Fig. 2 and 3), causing an information bottleneck at the deepest layers during forward propagation. This makes it easy for spatial information to be lost as the gradient flow reaches the deeper layers of the network, resulting in reduced data efficiency or the need for more complex network designs to achieve better performance.
To better evaluate information loss, the G-index is defined by the sum of two components: the absolute change in the maximum intensities and the absolute change in the minimum intensities of the feature maps. It quantifies the absolute difference in both the highest and lowest feature map intensities between two consecutive layers.
$GI=\sum^{L}_{l=1}(\left\|\Delta G_{min}^{l}\right\|+\left\|\Delta G_{max}^{l}\right\|)$ (4)
where $L$ is the depth of network. G-index, encapsulating the total of these absolute changes, acts as a measure directly proportional to the information loss within the network. (refer to Figure 3.)
3.3 Dense-Residual Group Auxiliary Supervision
Motivated by RRDB [43], Wang et al. suggested that incorporating dense-residual connections can aggregate multi-level feature information and stabilize training through residual learning [25, 34]. Nonetheless, CNN-based approaches, which primarily focus on learning local information, often overlook the importance of capturing long-range dependencies. Recent studies on SwinIR-based methods have concentrated on enlarging the receptive field [5, 6, 58] or enhancing the network’s capability to extract features [21, 44].
By adding dense connections within SwinIR-based blocks, we believe that it is possible to stabilize the gradient flow within each residual group during back-propagation, thereby addressing the issue of information bottlenecks and enhancing the receptive field. Consequently, this approach allows for achieving outstanding performance with simpler model architectures, or even by using shallower SISR networks.
@@ 4.1 Network Architecture
As shown in Fig. 4, DRCT is composed of three distinct components: shallow feature extraction, deep feature extraction and image reconstruction module.
@@
Shallow and deep feature extraction Given a low-resolution (LR) input $I_{LR}\in\mathbb{R}^{H\times W\times C_{in}}$ ($H$, $W$ and $C_{in}$ are the image height, width and input channel number, respectively), we use a $3\times 3$ convolutional layer $\mathcal{C}_{3}$()($\cdot$)( ⋅ ) [45] to extract shallow feature $F_{\text{0}}\in\mathbb{R}^{H\times W\times C}$ as
$F_{\text{0}}=\mathcal{C}_{3}(I_{LQ}),$ (5)
@@
Then, we extract deep feature which contains high-frequency spatial information $F_{DF}\in\mathbb{R}^{H\times W\times C}$ from $F_{\text{0}}$ and it can be defined as
$F_{DF}=H_{DF}(F_{0}),$ (6)
@@
where $H_{DF}(\cdot)$ is the deep feature extraction module and it contains $K$ Residual deep feature extraction group (RDFEG) and a convolutional layer $\mathcal{C}_{3}$. More specifically, intermediate features $F_{1},F_{2},\ldots,F_{K}$ and the output deep feature $F_{DF}$ are extracted block by block as
$F_{i}=RDFEG_{i}(F_{i-1}),\quad i=1,2,\ldots,K,$ (7)
$F_{DF}=\mathcal{C}_{3}(F_{K}),$ (8)
@@
Image reconstruction. We reconstruct the SR image $I_{SR}$ $\in\mathbb{R}^{H\times W\times C_{in}}$ by aggregating shallow and deep features, it can be defined as:
$I_{SR}=H_{REC}(F_{0}+F_{DF}),$ (9)
@@
where $H_{REC}(\cdot)$ is the function of the reconstruction module for fusing high-frequency deep feature $F_{DF}$ and low-frequency feature $F_{0}$ together to obtain SR result.
4.2 Residual Deep Feature Extraction Group and Swin-Dense-Residual-Connected Block
In developing the RDFEG, we take cues from [16] and [43], employing a residual-dense block (RDB) as the foundational unit. The reuse of feature maps emerges as the enhanced receptive field in the RDB’s feed-forward mechanism. The RDFEG contains a Swin-dense-residual-connected block (SDRCB) (refer to Figure 4), and the SDRCB can be defined as:
$X_{1}$ $=\text{PE}\left(\mathcal{C}_{\sigma}\left(\text{PUE}\left(\text{STL}(X)\right)\right)\right)$
$X_{2}$ $=\text{PE}\left(\mathcal{C}_{\sigma}\left(\text{PUE}\left(\text{STL}\left(\text{Cat.}(X,X_{1}),\right)\right)\right)\right)$
$X_{3}$ $=\text{PE}\left(\mathcal{C}_{\sigma}\left(\text{PUE}\left(\text{STL}\left(\text{Cat.}(X,X_{1},X_{2})\right)\right)\right)\right)$
$X_{4}$ $=\text{PE}\left(\mathcal{C}_{\sigma}\left(\text{PUE}\left(\text{STL}\left(\text{Cat.}(X,X_{1},X_{2},X_{3})\right)\right)\right)\right)$
$X_{5}$ $=\text{PE}\left(\mathcal{C}_{1}\left(\text{PUE}\left(\text{STL}\left(\text{Cat.}(X,X_{1},X_{2},X_{3},X_{4})\right)\right)\right)\right)$
$SDRCB(X)=\alpha\cdot X_{5}+X$ (10)
[…]
4.3 Same-task Progressive Training

### v4

@@ Abstract
In recent years, Vision Transformer-based approaches for low-level vision tasks have achieved widespread success. Unlike CNN-based models, Transformers are more adept at capturing long-range dependencies, enabling the reconstruction of images utilizing non-local information. In the domain of super-resolution, Swin-transformer-based models have become mainstream due to their capability of global spatial information modeling and their shifting-window attention mechanism that facilitates the interchange of information between different windows. Many researchers have enhanced model performance by expanding the receptive fields or designing meticulous networks, yielding commendable results. However, we observed that it is a general phenomenon for the feature map intensity to be abruptly suppressed to small values towards the network’s end. This implies an information bottleneck and a diminishment of spatial information, implicitly limiting the model’s potential. To address this, we propose the Dense-residual-connected Transformer (DRCT), aimed at mitigating the loss of spatial information and stabilizing the information flow through dense-residual connections between layers, thereby unleashing the model’s potential and saving the model away from information bottleneck. Experiment results indicate that our approach surpasses state-of-the-art methods on benchmark datasets and performs commendably at the NTIRE-2024 Image Super-Resolution (x4) Challenge. Our source code is available at https://github.com/ming053l/DRCT
@@ 1 Introduction
Figure 1: The feature map intensity on various benchmark datasets. We observed that feature map intensities decrease sharply at the end of SISR network, indicating potential information loss. In this paper, we propose DRCT to address this issue by enhancing receptive fields and adding dense-connections within residual blocks to mitigate information bottlenecks, thereby improving performance with a simpler model design.
@@
Following the foundational studies, CNN-based strategies [8, 35, 20, 19, 40, 39] have predominantly governed the super-resolution domain for an extended period. These strategies largely leverage techniques such as residual learning [61, 52, 46, 12, 36], or recursive learning [21, 18] for developing network architectures, significantly propelling the progress of super-resolution models forward.
CNN-based networks have achieved notable success in terms of performance. However, the inductive bias of CNN limits SISR models capture long-range dependencies. Their inherent limitations stem from the parameter-dependent scaling of the receptive field and the kernel size of convolution operator within different layers, which may neglect non-local spatial information within images.
To overcome the limitations associated with CNN-based networks, researchers have introduced Transformer-based SISR networks that leverage the capability to model long-range dependencies, thereby enhancing SISR performance. Notable examples include IPT [16] and EDT [33], which utilize pre-training on large-scale dataset like ImageNet [14] to fully leverage the capabilities of Vision Transformer [9] for achieving ideal SISR results. Afterwards, SwinIR [34] incorporates Swin-Transformer [26] into SISR, marked a significant advancement in SISR performance.
@@
While using Transformer-based SISR model for inference across various datasets, we observed a common phenomenon: the intensity distribution of the feature maps undergoes more substantial changes as the network depth increases. This indicates the spatial information and attention intensity learned by the model. However, there’s often sharp decrease towards the end of the network (refer to Figure 1) , shrinking to a smaller range. This phenomenon suggests that such abrupt changes might be accompanied by a loss of spatial information, indicating the presence of an information bottleneck.
Inspired by a series of works by Wang et al., such as the YOLO-family [50, 47], CSPNet [48], and ELAN [49], we consider that network architectures based on SwinIR, despite significantly enlarging the receptive fields through shift-window attention mechanism to address the small receptive fields in CNNs, are prone to gradient bottlenecks due to the loss of spatial information as network depth increases. This implicitly constrains the model’s performance and potential.
To address the issue of spatial information loss due to an increased number of network layers, we introduce the Dense-residual-connected Transformer (DRCT), designed to stabilize the forward-propagation process and prevent information bottlenecks. This is achieved by the proposed Swin-Dense-Residual-Connected Block (SDRCB), which incorporates Swin Transformer Layers and transition layers into each Residual Deep feature extraction Group (RDG). Consequently, this approach enhances the receptive field with fewer parameters and a simplified model architecture, thereby resulting in improved performance. The main contributions of this paper are summarised as follows:
  • We observed that as the network depth increases, the intensity of the feature map will gradually increase, then abruptly drop to a smaller range. This severe oscillation may be accompanied by information loss.
  • We propose DRCT by adding dense-connection within residual groups to stabilize the information flow for deep feature extraction during propagation, thereby saving the SISR model away from the information bottleneck.
  • By integrating dense connections into the Swin-Transformer-based SISR model, the proposed DRCT achieves state-of-the-art performance while maintaining efficiency. This approach showcases its potential and raises the upper-bound of the SISR task.
@@ 2.1 Vision Transformer-based Super-Resolution
IPT [16], a versatile model utilizing the Transformer encoder-decoder architecture, has shown efficacy in several low-level vision tasks. SwinIR [34], building on the Swin Transformer [26] encoder, employs self-attention within local windows during feature extraction for larger receptive fields and greater performance, compared to traditional CNN-based approaches. UFormer [53] introduces an innovative local-enhancement window Transformer block, utilizing a learnable multi-scale restoration modulator within the decoder to enhance the model’s ability to detect both local and global patterns. ART [64] incorporates an attention retractable module to expand its receptive field, thereby enhancing SISR performance. CAT [5] leverages rectangle-window self-attention for feature aggregation, achieving a broader receptive field. HAT [4] integrates channel attention mechanism [51] with overlapping cross-attention module, activating more pixels to reconstruct better SISR results, thereby setting new benchmarks in the field.
Figure 2: The feature map visualization displays, from top to bottom, SwinIR [34], HAT [4], and the proposed DRCT, with positions further to the right representing deeper layers within the network. For both SwinIR and HAT, the intensity of the feature maps is significant in the shallower layers but diminishes towards the network’s end. We consider this phenomenon implies the loss of spatial information, leading to the limitation and information bottleneck with SISR tasks. As for the proposed DRCT, the learned feature maps are gradually and stably enhanced without obvious oscillations. It represents the stability of the information flow during forward propagation, thereby yielding higher intensity in the final layer’s output. (zoom in to better observe the color-bar besides feature maps.)
@@ 2.2 Auxiliary Supervision and Feature Fusion
Auxiliary Supervision. Deep supervision is a commonly used auxiliary supervision method [31, 13] that involves training by adding prediction layers at the intermediate levels of the model [47, 48, 49]. This approach is particularly prevalent in architectures based on Transformers that incorporate multi-layer decoders. Another popular auxiliary supervision technique involves guiding the feature maps produced by the intermediate layers with relevant metadata to ensure they possess attributes beneficial to the target task [11, 28, 52, 61, 30]. Choosing the appropriate auxiliary supervision mechanism can accelerate the model’s convergence speed, while also enhancing its efficiency and performance.
Feature Fusion. Many studies have explored the integration of features across varying dimensions or multi-level features, such as FPN [22], to obtain richer representations for different tasks [36, 55]. In CNNs, attention mechanisms have been applied to both spatial and channel dimensions to improve feature representation; examples of which include RTCS [10] and SwinFusion [37]. In ViT [9], spatial self-attention is used to model the long-range dependencies between pixels. Additionally, some researchers have investigated the incorporation of channel attention within Transformers [62, 3] to effectively amalgamate spatial and channel information, thereby improving model performance.
@@ 3.1 Information Bottleneck Principle
$I(X,X)\geq I(Y,X)\geq I(Y,f_{\theta}(X))\geq I(X,g_{\phi}(f_{\theta}(X))),$ (1)
@@
In deep neural networks, $f_{\theta}(\cdot)$ and $g_{\phi}(\cdot)$ respectively represent the two consecutive layers in neural network. From equation (1), we consider that as the number of network layer becomes deeper, the information flow will be more likely to be lost. In term of SISR tasks, the general goal is to find the mapping function $F$ with optimized function parameters $\theta$ to maximize the mutual information between HR and SR image.
$F(\mathbf{I}_{LR};\theta)=\mathbf{I}_{SR};\\ \max_{\theta}I(\mathbf{I}_{HR};F(\mathbf{I}_{LR};\theta))$ (2)
@@ 3.2 Spatial Information Vanish in Super-resolution
Generally speaking, SISR methods [4, 5, 34, 58, 64, 6, 63, 32] can generally divided into three parts: (1) shallow feature extraction, (2) deep feature extraction, (3) image reconstruction. Among these methods, there is almost no difference between shallow feature extraction and image reconstruction. The former is composed of simple convolution layers, while the latter consists of convolution layers and upsampling layers. However, deep feature extraction differs significantly. Yet, their commonality lies in being composed of various residual blocks, which can be simply defined as:
$X^{l+1}=X^{l}+f_{\theta}^{l+1}(X^{l}),$ (3)
@@
Especially for SISR task, two methods of stabilizing information flow or training process are introduced:
Residual connection to learn local feature. Adopting residual learning allows the model to only update the differences between layers, rather than output the total information from a previous layer directly [12]. This reduces the difficulty of model training and prevents gradient vanishing locally [61]. However, according to our observations, while this design effectively transmits spatial information between different residual blocks, there may still be information loss.
Because the information within a residual block may not necessarily maintain spatial information, this ultimately leads to non-smoothness in terms of feature map intensity (refer to Fig. 2), causing an information bottleneck at the deepest layers during forward propagation. This makes it easy for spatial information to be lost as the gradient flow reaches the deeper layers of the network, resulting in reduced data efficiency or the need for more complex network designs to achieve better performance.
Dense connection to stabilize information flow. Incorporating dense connections into the Swin-Transformer based SISR model offers two significant advantages. Firstly, global auxiliary supervision. It effectively fuses the spatial information across different residual groups [52, 61], preserving high-frequency features throughout the deep feature extraction. Secondly, saving SISR model away from information bottleneck. By leveraging the integration of spatial information, the model ensures a smooth transmission of spatial information [46], thereby mitigating the information loss and enhancing the receptive field.
4 Motivation
Dense-Residual Group auxiliary supervision. Motivated by RRDB-Net [52], Wang et al. suggested that incorporating dense-residual connections can aggregate multi-level spatial information and stabilize the training process [35, 41]. We consider that it is possible to stabilize the information flow within each residual-groups during propagation, thereby saving SISR model away from the information bottleneck.
Dense connection with Shifting-window mechanism. Recent studies on SwinIR-based methods have concentrated on enlarging the receptive field [4, 5, 64, 6] by sophisticated WSA or enhancing the network’s capability to extract features [32, 53] for high-quality SR images. By adding dense-connections [15] within Swin-Transformer-based blocks [34, 26] in the SISR network for deep feature extraction, the proposed DRCT’s receptive field is enhanced while capturing long-range dependencies. Consequently, this approach allows for achieving outstanding performance with simpler model architectures [46], or even using shallower SISR networks.
@@ 5.1 Network Architecture
Figure 3: The overall architecture of the proposed DRCT and the structure of RDG and SDRCB.
@@
As shown in Figure 3, DRCT comprises three distinct components: shallow feature extraction, deep feature extraction, and image reconstruction module, respectively.
@@
Shallow and deep feature extraction. Given a low-resolution (LR) input $\mathbf{I}_{LR}\in\mathbb{R}^{H\times W\times C_{in}}$ ($H$, $W$ and $C_{in}$ are the image height, width and input channel number, respectively), we use a $3\times 3$ convolution layer Conv()($\cdot$)( ⋅ ) [54] to extract shallow feature $\mathbf{F}_{\text{0}}\in\mathbb{R}^{H\times W\times C}$ as
$\mathbf{F}_{\text{0}}=\text{Conv}(\mathbf{I}_{LQ}),$ (4)
@@
Then, we extract deep feature which contains high-frequency spatial information $\mathbf{F}_{DF}\in\mathbb{R}^{H\times W\times C}$ from $\mathbf{F}_{\text{0}}$ and it can be defined as
$\mathbf{F}_{DF}=H_{DF}(\mathbf{F}_{0}),$ (5)
@@
where $H_{DF}(\cdot)$ is the deep feature extraction module and it contains $K$ Residual Deep feature extraction Group (RDG) and single convolution layer $\text{Conv}(\cdot)$ for feature transition. More specifically, intermediate features $\mathbf{F}_{1},\mathbf{F}_{2},\ldots,\mathbf{F}_{K}$ and the output deep feature $\mathbf{F}_{DF}$ are extracted block by block as
$\mathbf{F}_{i}=\text{RDG}_{i}(\mathbf{F}_{i-1}),\quad i=1,2,\ldots,K,$ (6)
$\mathbf{F}_{DF}=\text{Conv}(\mathbf{F}_{K}),$ (7)
@@
Image reconstruction. We reconstruct the SR image $\mathbf{I}_{SR}$ $\in\mathbb{R}^{H\times W\times C_{in}}$ by aggregating shallow and deep features, it can be defined as:
$\mathbf{I}_{SR}=H_{\text{rec}}(\mathbf{F}_{0}+\mathbf{F}_{DF}),$ (8)
@@
where $H_{\text{rec}}(\cdot)$ is the function of the reconstruction for fusing high-frequency deep feature $\mathbf{F}_{DF}$ and low-frequency feature $\mathbf{F}_{0}$ together to obtain SR result.
5.2 Deep Feature Extraction
Residual Deep feature extraction Group. In developing of RDG, we take cues from RRDB-Net [52] and RDN [61], employing a residual-dense block (RDB) as the foundational unit for SISR. The reuse of feature maps emerges as the enhanced receptive field in the RDG’s feed-forward mechanism. To expound further, RDG with several SDRCB enhances the capability to integrate information across different scales, thus allowing for a more comprehensive feature extraction. RDG facilitates the information flow within residual group, capturing the local features and spatial information group by group.
Swin-Dense-Residual-Connected Block. In purpose of capturing the long-range dependency, we utilize the shifting window self-attention mechanism of Swin-Transformer Layer (STL) [26, 34] for obatining adaptive receptive fields, complementing RRDB-Net by focusing on multi-level spatial information. This synergy leverages STL to dynamically adjust the focus of the model based on the global content of the input, allowing for a more targeted and efficient extraction of features. This mechanism ensures that even as the depth of the network increases, global details are preserved, effectively enlarging and enhancing the receptive field without compromising. By integrating STLs with dense-residual connections, the architecture benefits from both a vast receptive field and the capability to hone in on the most relevant information, thereby enhancing the model’s performance in SISR tasks requiring detailed and context-aware processing. For the input feature maps $\mathbf{Z}$ within RDG, the SDRCB can be defined as:
$\mathbf{Z}_{j}=H_{\text{trans}}(\text{STL}([\mathbf{Z},...,\mathbf{Z}_{j-1}])),j=1,2,3,4,5,$ (9)
$\text{SDRCB}(\mathbf{Z})=\alpha\cdot\mathbf{Z}_{5}+\mathbf{Z},$ (10)
where []delimited-[][$\cdot$][ ⋅ ] denotes the concatenation of multi-level feature maps produced by the previous layers. $H_{\text{trans}}(\cdot)$ refers to the convolution layer with a LeakyReLU activate function for feature transition. The negative slope of LeakyReLU is set to $0.2$. $\text{Conv}_{1}$ is the $1\times 1$ convolution layer, which is used to adaptively fuse a range of features with different levels [42]. $\alpha$ represents residual scaling factor, which is set to $0.2$ for stabilizing the training process [52].
5.3 Same-task Progressive Training Strategy